### v1

@@ 4.3 Caveats
Here discuss a few limitations in our model. First, for simplicity we treat $\tau_{\rm a}$ and $\tau_{\rm e}$ as free but constant parameters and two planets have the same eccentricity damping timescales. However, in reality $\tau_{\rm a}$ and $\tau_{\rm e}$ depend on the gas disk mass, planet mass and orbital distance. For instance, $\tau_{\rm a}$ and $\tau_{\rm e}$ decrease as planet mass increases, which scale with $\mu^{-1}$. The two trapping criteria exhibit steeper decreasing functions of $\mu$ such that $\tau_{\rm a,\ \text{slow\ mig}}{\propto}\mu^{-4/3}$, and $(\tau_{\rm a}\tau_{\rm e})_{\rm\text{weak\ damp}}{\propto}\mu^{-2}$. This suggests that planets become more easily trapped as their mass increases. On the other hand, since the planets can grow their masses at different rates, the planet mass ratio could be time varied. For instance, if the inner planet is initially more massive than the outer planet ($q{\gg}1$), the resonance trapping is the stable regime. When the outer planet grows and the mass ratio decreases, the system may transition into an overstable trap or even escape. Moreover, both hydrodynamic simulations and analytical work have shown that the migration and eccentricity damping rates differ from those in the near circular case when planets attain high eccentricities (Cresswell & Nelson, 2008; Ida et al., 2020). Future studies are required to investigate how these factors influence our analysis.
@@
Second, we only consider a static disk. The configuration may change when disk dissipation is taken into account. In this case, the migration slows down as the gas density drops, making the resonance trapping easier. However, for stability/overstability and escape criteria, the situation is less straightforward. This is because $\tau_{\rm a}/\tau_{\rm e}$ depends on the aspect ratio of the disk but not the surface density. The disk thermal structure relies on the heating mechanism and therefore a specific disk model. As such, our framework cannot predict whether the planet pairs in the stable or overstable regimes will remain trapped as gas disk dissipates.
In addition, the disk is assumed to be laminar here. In a more realistic turbulent disk, density fluctuations generated by turbulence can result in additional stochastic torques (Goldberg & Batygin, 2023; Rein, 2012). This effect cannot be captured by our simplified approach that assumes static migration and eccentricity damping timescales. However, the outcomes of resonance trapping may be altered by considering the stochastic torques, which is a key focus of our future work.
@@ 5 Conclusion
We have built an analytical framework suited to study the resonant trapping, stability, and escape conditions for a two-planet pair that undergoes the convergent type I migration in a gaseous disk. The $2$:$1$ first order mean motion resonance is specifically investigated and the key outcomes are illustrated in Fig 6.
Planets can be trapped in resonances under the circumstances of slow migration and weak damping (Eqs. (8) and (11)). When the mass of inner planet is comparable or less massive than the outer planet ($q{\lesssim}1$), after capture the stability weakens as $\tau_{\rm a}/\tau_{\rm e}$ decreases. The system transitions from a stable trap to an overstable trap (Eq. 22), and finally to escape (Eq. 23). This corresponds to the resonant states changing from statically stable to dynamically stable, and ultimately to dynamically unstable. In contrast, when the mass of inner planet is much more massive than the outer planet ($q{\gg}1$), the traps remain permanently stable.
We provide the general analytical criteria for planets with an arbitrary mass ratio $q$ in Sec. 2. The corresponding formulas are also summarized as follows.
$\begin{cases}{\displaystyle(\tau_{e}\tau_{a})_{\rm weak,e}=\frac{1}{2(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vspace{0.1cm}\\ {\displaystyle(\tau_{a})_{\rm slow,mig}=\frac{1}{2[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vspace{0.1cm}\\ {\displaystyle\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (33)
where
$h(q)$ $=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (34)
We validate the theoretical predictions with numerical N-body simulations (Figures 8, 8 and 9). Furthermore, within the framework, we show that resonance trapping cannot occur when planets migrate divergently. We also explored the dependence of the semi-major axis and the stellar mass based on minimum mass solar nebular (MMSN) model, which indicates planets are more likely to be resonance-trapped with increasing semi-major axis and decreasing stellar mass.
@@ Acknowledgements.
This work is supported by REBOUND (Rein & Liu, 2012) and REBOUNDX (Tamayo et al., 2020)
@@
where the contribution from $\dot{\varpi_{\rm i}}$ and $\dot{\varpi_{\rm o}}$ are neglected. With no additional effects from the gaseous disk such as orbital decay and eccentricity damping, Lagrange’s equations can be written as
@@
$\dot{e_{\rm i}}f_{\rm d,i}-\dot{e_{\rm o}}f_{\rm d,o}^{{}^{\prime}}=-(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})\sin{\varphi},$ (A4)
@@
$d(e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{d,o^{{}^{\prime}}}){=}\pm\frac{(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}})}{\omega_{0}}\sin{\frac{\varphi}{2}}d{\varphi}.$ (A8)
@@
The maximum eccentricity occurs when $\varphi=0$, which provides the eccentricity described in Sec. 2.1.1.
@@
$\begin{cases}{\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}]^{2/3}\frac{e_{\rm i}^{2}}{4},\ \ \hfill[\mbox{$q\ll 1$}],\vspace{0.1cm}}\\ {\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm i}}\right)^{2/3}[\frac{j^{2}}{f_{\rm d,o}^{{}^{\prime}}}]^{2/3}\frac{e_{\rm o}^{2}}{4},\ \ \hfill[\mbox{$q\gg 1$}].}\end{cases}$ (B2)
@@
where $e^{2}=e_{\rm i}^{2}+e_{\rm o}^{2}$. When $\eta_{\rm eq}{<}-3/2$ (equivalently $\Phi_{\rm eq}>1$), a separatrix point occurs and the system can cross this separatrix and escape from resonance (see Sec. 2.2.2). By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ (Eq. 4) into Eq. (B3), we obtain the general escape criterion as
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm esc}=\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}.$ (B4)
@@
$h(q)$ $=$ $\tau_{a}/\tau_{e}\left(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2}\right)$ (B5)
$=$ $\frac{1}{2}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$
@@
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm overstab}=\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q),$ (B6)
@@
B.3: Dependence on the mass ratio $q$
We run simulations to testify to our proposed overstability and escape criteria. Figure B.1 shows the numerical results of $\tau_{\rm a}/\tau_{\rm e}$ as a function of $q$. We fix $m_{\rm i}{=}10\ M_{\oplus}$ and set $\tau_{\rm a}{=}3\ (\tau_{\rm a})_{\rm slow,mig}$ to fulfill the trapping criteria. The solid lines denote the analytical criteria, while the colored dots represent the simulation results.
@@
The numerical results agree pretty well with the analytical expectation at $q{\leq}1$. The simulations depict a transition from stable trap (green dots) to overstable trap (yellow dots), and eventually enter escape (purple dots) regimes as $\tau_{\rm a}/\tau_{\rm e}$ decreases, and our analytical criteria separate these three regimes. In contrast, when $q\geq 1$, there are still two points in Figure B.1 where the simulations do not fully match the analysis.
@@
First, the stability/overstability and escape criteria exhibit the opposite dependence $q$, with the two lines intersecting at $q{\sim}2$. For $q{\geq}2$, one would expect that the systems reside region below the overstability criterion (yellow line) correspond to escape. However, simulations show that all systems in this region are trapped overstably (no escape occurs). We attribute this to the fact that when $q{\geq}2$, the angular momentum deficit (AMD) (Laskar 1997) of the inner planet is larger than that of the outer planet, limiting the amplitude of oscillations in overstability (see Appendix B.4 for detailed derivation). As a result, the perturbations cannot grow sufficiently to cause escape from resonance. The dashed purple line in Figure B.1 represents the AMD criterion when the second-order terms of the eccentricities are considered. This new criterion is in good agreement with our numerical simulations.
Second, our overstable criterion suggests $\tau_{\rm a}/\tau_{\rm e}{<}0$ at $q{\geq}3$, indicating that all traps should be stable in this regime. However, in Figure B.1, it is evident that near $q\sim 4$ and $\tau_{a}/\tau_{e}\sim 10^{2}$, the planets remain in an overstability state. This is because we only considered the first-order term of the disturbing function. Second-order terms are neglected here, weakening the inner planet effect on the outer planet and leading to an overstable trap.
Figure B.1: Simulation results of $2:1$ MMR stability in $q-\tau_{a}/\tau_{e}$ parameter space with the mass of the inner planet $m_{\rm i}=10\ \times M_{\oplus}$. Green points indicate stable resonance capture, yellow points indicate overstable resonance capture, and purple points indicate the planets eventually escape from the resonance. The lines represents our analytical criteria. It can be observed that our criteria overall align well with the numerical results.
@@
We expand the disturbing functions to the second order in $e_{\rm i}$ and $e_{\rm o}$, which gives
@@
$AMD\sim\mu\sqrt{a}(1-\sqrt{1-e^{2}})\sim\mu\sqrt{a}e^{2}.$ (B14)
@@
is satisfied. Under this condition, the amplitude of oscillations following overstability is limited, and therefore the perturbations cannot grow sufficiently to cause escape. By substituting Eqs. (B12) and (B13) to eliminate either $e_{\rm i}$ and $e_{\rm o}$ under the condition $\dot{\varpi_{\rm i}}=\dot{\varpi_{\rm o}}$, we obtain the relationship between $e_{\rm i,o}$ and $q$. Assuming $e_{\rm i}$ takes the form: $e_{\rm i}^{2}={\tau_{e}}/{2j\tau_{a}}$, and substituting this into the derived form, we obtain our AMD criterion as
$\left(\frac{\tau_{a}}{\tau_{e}}\right)_{\rm AMD}=\frac{2}{j}\left[\frac{(f_{3}-K_{1})+q^{{}^{\prime}2}(f_{5}+K_{1})}{q^{{}^{\prime}}f_{\rm d,o}^{{}^{\prime}}+f_{\rm d,i}}\right]^{2},$ (B16)
where $q^{{}^{\prime}2}\equiv q\sqrt{\alpha}$. The AMD criterion aligns well with the numerical results (Figure B.1).

### v2

@@ 4 Discussion
In this section, we outline a few limitations in our model. First, for simplicity we treat $\tau_{\rm m}$ and $\tau_{e}$ as free but constant parameters and two planets have the same eccentricity damping timescales. However, in reality $\tau_{\rm m}$ and $\tau_{e}$ depend on the gas disk mass, the planet mass, and orbital distance. For instance, in the linear type I regime, both $\tau_{\rm m}$ and $\tau_{e}{\propto}\mu^{-1}$, while the two trapping criteria exhibit steeper decreasing functions of $\mu$ such that $\tau_{\rm m,\ \text{slow\ mig}}{\propto}\mu^{-4/3}$, and $(\tau_{\rm m}\tau_{e})_{\rm\text{weak\ damp}}{\propto}\mu^{-2}$. This suggests that planets become more easily trapped as their masses increase. Nevertheless, we note that the relation between planet mass and trapping outcome behavors more complicated. When the planets’ masses further increase and exceed the gap-opening threshold, they gradually switch to Type-II migration (Lin & Papaloizou, 1986; Kanagawa et al., 2018). In this regime, the planet’s mass influences migration and eccentricity damping in a non-trivial way (Kanagawa et al., 2018; Kajtazi et al., 2023; Pichierri et al., 2023, 2024). This additional complexity is not accounted for in our current analysis. On the other hand, since the planets can grow their masses at different rates, the planet mass ratio could change over time. For instance, if the inner planet is initially more massive than the outer planet ($q{\gg}1$), the resonance trapping is in the stable regime. When the outer planet grows and the mass ratio decreases, the system may transition into an overstable trap or even escape. Future studies should explore how these interactions modify the trapping criteria and how the evolving planet mass ratio affects the long-term dynamics of the system.
Furthermore, both hydrodynamic simulations and analytical work have shown that the migration and eccentricity damping rates differ from those in the near circular case when planets attain high eccentricities (Cresswell & Nelson, 2008; Bitsch & Kley, 2011; Bitsch et al., 2013; Xu et al., 2018; Ida et al., 2020; Pichierri et al., 2023, 2024; Fairbairn & Rafikov, 2025). For instance, Xu et al. (2018) showed that nonlinear eccentricity damping can lead to higher equilibrium eccentricities, enhance the stability of captured resonances.
In addition, our analysis carried out so far assumes a laminar disk. In a more realistic turbulent disk, turbulence-induced density fluctuations can generate additional stochastic torques (Laughlin et al., 2004; Ogihara et al., 2006; Rein, 2012; Okuzumi & Ormel, 2013; Hühn et al., 2021; Wu et al., 2024). These effects are not accounted for in the aforementioned simplified framework. Such stochastic torques may considerably alter the resonance trapping outcomes (Batygin & Adams, 2017; Goldberg & Batygin, 2023; Chen et al., 2025).
@@
Lastly, we only consider a static disk without considering the gas density decay due to disk dissipation. As disk gas disperses, the planet migration slows down and resonance trapping is easier to occur. However, for stability/overstability and escape criteria, the situation is less straightforward. This is because $\tau_{\rm m}/\tau_{e}$ depends on the aspect ratio of the disk but not the surface density. The disk thermal structure relies on the heating mechanism. While our framework remains applicable and can predict the evolution of resonance configurations as long as $\tau_{m}$ and $\tau_{e}$ are specified, understanding how these timescales change with disk dissipation is essential for making such predictions. However, this aspect is not within the scope of our current study.
@@ 5 Conclusion
We have constructed an analytical framework suited to study the resonant trapping, stability, and escape conditions for two-planet systems that undergo convergent type I migration in the protoplanetary disks. The planet-disk interaction is assumed to be parameterized by two key quantities: angular momentum damping and eccentricity damping timescales $\tau_{\rm m}$ and $\tau_{e}$. We present the analytical criteria for general $j$:j1j-1 resonances and for planet pairs with arbitrary mass ratios. The outcomes can be classified into four regimes: direct resonance crossing, stable resonance trap, overstable resonance trap and escape after temporary trap. The proposed criteria and regimes, which are validated through numerical N-body simulations, can be illustrated in a $\tau_{\rm m}/\tau_{\rm e}$-$\tau_{e}$ plot (Figure 6).
We find that the planets can be trapped in resonances under slow migration and weak damping (Eqs. 8 and 11). When the mass of the inner planet is comparable to or less massive than the outer planet ($q{\lesssim}1$), after capture, the stability weakens as $\tau_{\rm m}/\tau_{e}$ decreases. The system transitions from a stable trap to an overstable trap (Eq. 22), and finally to escape (Eq. 23). This corresponds to the resonant states changing from statically stable to dynamically stable, and ultimately to dynamically unstable. In contrast, when the mass of the inner planet is much more massive than the outer planet ($q{\gg}1$), the traps remain permanently stable. The key formulas of different regimes are presented in Eqs. (13) and (24). Future studies are needed to connect this theoretical framework with observations (Dai et al., 2024; Hu et al., 2025).
@@ Acknowledgements.
The authors appreciate the constructive comments from the referee. BL is supported by the National Key R&D Program of China (2024YFA1611803), the National Natural Science Foundation of China (Nos. 12222303 and 12173035), and the start-up grant of the Bairen program from Zhejiang University. The simulations and analysis presented in this article were carried out on the SilkRiver Supercomputer of Zhejiang University. Software: REBOUND (Rein & Liu, 2012) and REBOUNDX (Tamayo et al., 2020).
@@
where the contribution from $\dot{\varpi_{\rm i}}$ and $\dot{\varpi_{\rm o}}$ are neglected. With no additional effects from the gaseous disk such as angular momentum damping and eccentricity damping, Lagrange’s equations can be written as
@@
$\dot{e_{\rm i}}f_{\rm d,i}-\dot{e_{\rm o}}f_{\rm d,o}^{{}^{\prime}}=-(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})\sin{\varphi},$ (A4)
@@
$d(e_{\rm i}f_{\rm d,i}-e_{\rm o}f_{d,o^{{}^{\prime}}}){=}\pm\frac{(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})}{\omega_{0}}\sin{\frac{\varphi}{2}}d{\varphi}.$ (A8)
@@
The maximum eccentricity occurs when $\varphi=0$, which provides the eccentricity described in Section 2.1.1.
@@
$\begin{cases}{\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm o}}\right)^{2/3}[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}]^{2/3}\frac{e_{\rm i}^{2}}{4},\ \ \hfill[\mbox{$q\ll 1$}],\vskip 2.84544pt}\\ {\displaystyle\Phi\sim\left(\frac{3}{\mu_{\rm i}}\right)^{2/3}[\frac{j^{2}}{f_{\rm d,o}^{{}^{\prime}}}]^{2/3}\frac{e_{\rm o}^{2}}{4},\ \ \hfill[\mbox{$q\gg 1$}].}\end{cases}$ (B2)
@@
where $e^{2}=e_{\rm i}^{2}+e_{\rm o}^{2}$. When $\eta_{\rm eq}{<}-3/2$ (equivalently $\Phi_{\rm eq}>1$), a separatrix point occurs and the system can cross this separatrix and escape from resonance (see Section 2.2.2). By substituting $e_{\rm i,eq}$ and $e_{\rm o,eq}$ (Eq. 4) into Eq. (B3), we obtain the general escape criterion as
$\frac{\tau_{\rm m}}{\tau_{e}}<\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}.$ (B4)
@@
$h(q)$ $=$ $\tau_{\rm m}/\tau_{e}\left(e_{\rm i,eq}^{2}+e_{\rm o,eq}^{2}\right)$ (B5)
$=$ $\frac{1}{(1+q\sqrt{\alpha})}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$
@@
$\frac{\tau_{\rm m}}{\tau_{e}}>\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q),$ (B6)
@@
B.3: Dependence on planet mass ratio
We run simulations to test our proposed overstability and escape criteria with different planet mass ratios by fixing $m_{\rm i}{=}10\ M_{\oplus}$. We set $\tau_{\rm m}{=}3\ (\tau_{\rm m})_{\rm slow,mig}$ to fulfill the trapping criteria and vary $\tau_{e}$. The result is presented in Figure B.1, where the solid and dashed lines denote the analytical criteria from ours and from Deck & Batygin (2015).
@@
The simulations agree pretty well with the analytical expectation at $q{\leq}1$, with our criteria slightly better than Deck & Batygin (2015). The simulations depict a transition from stable trap (green squares) to overstable trap (yellow circles), and eventually become escape (purple triangles) as $\tau_{\rm m}/\tau_{e}$ decreases, and our analytical criteria separate these three regimes. However, at $q{\geq}1$, there are still two mismatches between simulations and analysis in Figure B.1.
@@
First, the stability/overstability and escape criteria exhibit an opposite trend in q, with the two lines intersecting at $q{\sim}2$. For $q{\gtrsim}2$, one would expect that the systems in the region below the overstability criterion (yellow line) should escape. However, simulations show that all systems in this region are trapped overstably (no escape occurs). We attribute this to the fact that when $q{\gtrsim}2$, the angular momentum deficit (AMD) (Laskar 1997) of the inner planet is higher than that of the outer planet. The exchange of angular momentum between these two planets weakens the oscillations of the amplitude of the outer planet’s eccentricity (see Appendix B.4 for detailed derivation). Thus, the perturbations cannot grow sufficiently to cause escape, and the system is refined in the overstable trap regime. The solid blue line in Figure B.1 represents the AMD criterion up to the second-order eccentricity terms. For comparison, the blue dotted line is the AMD criterion with only the first-order eccentricity terms. We find that the the AMD criterion with the second-order eccentricity terms separates the overstable trap and escape regimes at $q{\gtrsim}2$, which is in good agreement with our simulations.
Second, our overstable criterion suggests at $q{\geq}3$, all traps are stable. However, in Figure B.1, it is evident that near $q\sim 4$ and $\tau_{\rm m}/\tau_{e}\sim 2\times 10^{2}$, the planets remain in an overstable state. We suspect the reason might be that we only considered the first-order term of the disturbing function. When the second-order term is included, the angular momentum transfer from the inner planet to the outer one becomes weaker. This leads to the shift of the regimes between stable and overstable to a higher $q$. A detailed analytical investigation for the offset seen in Fig. B.1 is nevertheless beyond the scope of this work.
Besides, we also extend the test to the 33:22 MMR in Fig. B.2. We also find that our analysis agrees well with the numerical simulations.
Figure B.1: Simulations of 22:11 MMR capture and stability in $q-\tau_{\rm m}/\tau_{e}$ parameter space. The mass of the inner planet $m_{\rm i}=10\ M_{\oplus}$. Green: stable resonance capture. Yellow: overstable resonant capture. Purple: resonance escape after temporary capture. The solid lines are the criteria derived in this study, whereas the dashed lines represent the stability and escape criteria of Deck & Batygin (2015) (their Eqs. 15 and 21). The blue solid and dotted lines represent our second-order (Eq. B16) and first-order (Eq. B17) AMD criteria, respectively.
Figure B.2: Same as Fig B.1, but for the 33:22 MMR.
@@
We expand the disturbing functions up to the second order in $e_{\rm i}$ and $e_{\rm o}$, which gives
@@
$AMD_{\rm i,o}\sim\mu_{i,o}\sqrt{a_{\rm i,o}}(1-\sqrt{1-e_{\rm i,o}^{2}})\sim\mu_{\rm i,o}\sqrt{a_{\rm i,o}}e_{\rm i,o}^{2}.$ (B14)
@@
is satisfied. Under this condition, the amplitude of oscillations following overstability is limited, and therefore the perturbations cannot grow sufficiently to cause escape. By substituting Eqs. (B12) and (B13) to eliminate either $e_{\rm i}$ and $e_{\rm o}$ under the condition $\dot{\varpi_{\rm i}}=\dot{\varpi_{\rm o}}$, we obtain the relationship between $e_{\rm i,o}$ and $q$. Assuming $e_{\rm i}$ takes the form: $e_{\rm i}^{2}={\tau_{e}}/{j\tau_{\rm m}}$, and substituting this into the derived form, we obtain our second-order AMD criterion as
$\left(\frac{\tau_{\rm m}}{\tau_{e}}\right)<\frac{1}{j}\left[\frac{(f_{3}-K_{1})+q^{{}^{\prime}2}(f_{5}+K_{1})}{q^{{}^{\prime}}f_{\rm d,o}^{{}^{\prime}}+f_{\rm d,i}}\right]^{2},$ (B16)
where $q^{{}^{\prime}2}\equiv q\sqrt{\alpha}$. We can also obtain the first-order AMD criterion by substituting Eq. (5) and (4b) into Eq. (B15):
$q>\left(\frac{f_{\rm d,i}}{f_{\rm d,o}^{{}^{\prime}}}\right)^{2}\frac{1}{\sqrt{\alpha}}.$ (B17)
By comparison, the inclusion of the second-order terms induce a $\tau_{\rm m}/\tau_{e}$ dependency, which is absent in the first-order approximation. We plot our AMD criteria in Figs. B.1 and B.2 and find that the second-order AMD criterion exhibits improved agreement with the numerical results.
C: Extension to Unequal Eccentricity Damping Timescales
We extend the framework to the planet pair systems with non-equal eccentricity damping. Equation (3) can be written as
$-[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}](e_{\rm i}f_{\rm d,i}\sin{\varphi_{\rm i}}+e_{\rm o}f_{\rm d,o}^{{}^{\prime}}\sin{\varphi_{\rm o}})$ (B18)
$=\frac{1}{\tau_{\rm m}}+\frac{e_{\rm o}^{2}}{\tau_{e\rm,o}}-\frac{e_{\rm i}^{2}}{\tau_{e\rm,i}}.$
The equilibrium eccentricities read
$e_{\rm i,eq}^{2}$ $=$ $\frac{\tau_{e,\rm i}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,i}^{2}}{jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e\rm,i}}{\tau_{e\rm,o}}f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (B19a)
$e_{\rm o,eq}^{2}$ $=$ $\frac{\tau_{e,\rm i}}{\tau_{\rm m}}\frac{1}{1+q\sqrt{\alpha}}\frac{f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e\rm,i}}{\tau_{e\rm,o}}f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}}.$ (B19b)
The key criteria can be re-formulated as
$\begin{cases}{\displaystyle\tau_{e,\rm i}\tau_{\rm m}>\frac{1}{(1+q\sqrt{\alpha})}\frac{1}{[jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e,\rm i}}{\tau_{e,\rm o}}f_{\rm d,o}^{{}^{\prime},2}q\sqrt{\alpha}]}\frac{1}{(\mu_{\rm o}n_{\rm i}\alpha)^{2}}}\\ {\displaystyle}\hfill[\mbox{weak e-damping}],\vskip 2.84544pt\\ {\displaystyle\tau_{\rm m}>\frac{1}{[(j-1)\mu_{\rm o}n_{\rm i}\alpha+j\mu_{\rm i}n_{\rm o}]}\frac{[(j-1)^{2}\mu_{\rm o}n_{\rm i}^{2}\alpha+j^{2}\mu_{\rm i}n_{\rm o}^{2}]^{1/3}}{3^{1/3}(\mu_{\rm o}n_{\rm i}\alpha f_{\rm d,i}^{2}+\mu_{\rm i}n_{\rm o}f_{\rm d,o}^{{}^{\prime}2})^{2/3}}}\\ {\displaystyle}\hfill[\mbox{slow migration}],\vskip 2.84544pt\\ {\displaystyle\frac{\tau_{\rm m}}{\tau_{e,\rm i}}>\left(\frac{3}{\mu_{o}}\right)^{2/3}h(q)\ \ \left[\frac{(j-1)^{2}}{f_{\rm d,i}\alpha}\right]^{2/3}(j-1)^{2/3}f(q)}\\ {\displaystyle}\hfill[\mbox{stability criterion}],\vskip 2.84544pt\\ {\displaystyle\frac{\tau_{\rm m}}{\tau_{e,\rm i}}<\left(\frac{3}{\mu_{o}}\right)^{2/3}\frac{h(q)}{4}\ \ \left[\frac{(j-1)^{2}+j^{2}q}{f_{\rm d,i}\alpha+f_{\rm d,o}^{{}^{\prime}}q^{2}}\right]^{2/3}}\\ {\displaystyle}\hfill[\mbox{escape criterion}],\end{cases}$ (B20)
where
$h(q)$ $=$ $\frac{1}{(1+q\sqrt{\alpha})}\frac{f_{\rm d,i}^{2}+f_{\rm d,o}^{{}^{\prime}2}q^{2}\alpha}{jf_{\rm d,i}^{2}+(j-1)\frac{\tau_{e,\rm i}}{\tau_{e,\rm o}}f_{\rm d,o}^{{}^{\prime}2}q\sqrt{\alpha}},$ (B21)
$f(q)$ $=$ $1-\left(\frac{f_{\rm d,o}^{{}^{\prime}}}{f_{\rm d,i}}\right)^{2}q^{2}\alpha.$ (B22)
These criteria have also been validated and compared with those of Deck & Batygin (2015) through simulations by incorporating different eccentricity dampings of two planets at $\tau_{e,\rm i}/\tau_{e,\rm o}{=}2$ (see Fig. C.1). However, we also find that the analytical formulas increasingly deviate from the simulation results as the ratio $\tau_{e,i}/\tau_{e,o}$ increases, particularly for values greater than 10. We suspect that this is because, in our derivation, $\cos{\varphi_{\rm i}}{=}\cos{\varphi_{\rm o}}{=}1$ is assumed, whereas it can not hold for extreme eccentricity damping ratios. A detailed investigation is nevertheless beyond the scope of this work.
Figure C.1: Same as Fig 9, but $\tau_{e,\rm i}/\tau_{e,\rm o}=2$.